OpenStudy (anonymous):

Consider the endpoint problem y''+ λy = 0; y(−1) = 0, y'(1) = 0. Find the positive eigenvalues and associated eigenfunctions.

OpenStudy (anonymous):

idk what that is sorry

OpenStudy (anonymous):

The characteristic equation is \(r^2 +\lambda =0\) hence \(r_{1,2}= \pm\sqrt{-\lambda}\) hence \(r_{1,2}=\pm i\sqrt{\lambda}\)

OpenStudy (anonymous):

We have the solution for the ODE is \(y=C_1cos(\sqrt{\lambda} x)+C_2 sin(\sqrt{\lambda}x)\)

OpenStudy (anonymous):

Replace y(-1) =0 to get the first equation Take derivative of y and replace y'(1) =0 to get the second equation Then solve for \(C_1,C_2\) I think you can handle it, right?